OpenStudy (anonymous):

solve for x? logbase2(5-x)-logbase2(2+x^2)=1?

OpenStudy (anonymous):

but...howd you get there?

OpenStudy (anonymous):

\[\log_{2} (5-x)-\log_{2}(2+x^2)=1\]\[\log_{2} (5-x)/(2+x^2)=\log_{2}2\]\[(5-x)/(2+x^2)=2\]\[5-x=4+2x^2\]\[2x^2+x-1=0\] \[Δ=1^2-4\times(-1)2=9\]\[x_1,_2=(-1\pm \sqrt{9})/2\times2\] \[x=1/2\]or\[x=-1\]

OpenStudy (anonymous):

excellent work...........i just guessed :)

OpenStudy (luffingsails):

Hi Christal, The key in what Johnny wrote is recognizing that the right side of the equation can be rewritten: \[\log_{2} 2 = 1\] After that, every term in your equation has log base 2... so then you can effectively eliminate them after some grouping. remember the log rules earlier? Starting with: \[\log_{2} (5-x)-\log_{2}(2+x^{2}) = 1 = \log_2{2}\] Remembering that (log a - log b = log a/b), we can get \[\log_{2} \frac{5-x}{2+x^{2}}=\log_{2}2\] Now, since they are all log base 2, get rid of that stupid log! \[\frac{(5-2)}{2+x^{2}}=2\] From there simplify and solve!

OpenStudy (anonymous):

wow thanks that really helped!

OpenStudy (anonymous):

could uu help me with another?

OpenStudy (luffingsails):

Hi Kmousou. I just learned today. if you have the fraction x/y, in the equation editor just type frac{x}{y}

OpenStudy (luffingsails):

Sorry, site seems really slow for me, Christal. Sure, I will help if I can.

OpenStudy (anonymous):

okay. h ow about ln(2x+7)=-1?

OpenStudy (luffingsails):

Give you a hint first, remember: \[\ln e^{x} = x\] So, your right side equation (the number -1) can be written as: \[\ln e^{-1}\] From there you can effectively get rid of your ln. Make sense so far? So what would you do next?

OpenStudy (anonymous):

ln2x times ln7= lne^-1..so get rid of ln?

OpenStudy (luffingsails):

Actually, the left hand side is fine already. It is in a good form for solving the equation. The goal is to have a single ln(of some values) = ln(of some values). So, \[\ln (2x+7) = \ln e^{-1}\] then, \[2x+7 = e^{-1} = \frac{1}{e}\] Then solve for x.

OpenStudy (luffingsails):

Sorry, \[\ln e^{1}\] Should be \[\ln e^{-1}\]

OpenStudy (luffingsails):

lol... should read (ln of e to the -1)... not sure why the -1 isn't showing up.

OpenStudy (anonymous):

its alright i got it! yayyy! except the next one...

OpenStudy (anonymous):

i see the negative one

OpenStudy (luffingsails):

!! cool.

OpenStudy (anonymous):

[ln(x)]^2-2ln(x)=0. is there a difference?

OpenStudy (luffingsails):

So, if I understand what you wrote: \[(\ln x)^{2} - 2 \ln {x} = 0\] Remember, the goal is to put a ln(something) = ln (something), so \[(\ln x)^{2} = 2 \ln {x}\] Ahhgg, going to run out of time to do this one right, but ln(x) * ln(x) = ln(x+x) (from the rules!) = ln(2x) Look up the power rule for logs because 2 * ln(x) can be rewritten too. Once you have done that, you have the form you need to simplify and drop the ln.

OpenStudy (anonymous):

thaaaank youu!! so much

OpenStudy (luffingsails):

you're welcome. See you later.